# Exploring the 6G Potentials: Immersive, Hyper Reliable, and Low-Latency Communication

### v1

@@ Abstract
The transition towards the sixth-generation (6G) wireless telecommunications networks introduces significant challenges for researchers and industry stakeholders. The 6G technology aims to enhance existing usage scenarios, particularly supporting innovative applications requiring stringent performance metrics. Among the key performance indicators (KPIs) for 6G, immersive throughput, hyper-reliability, and hyper-low latency must be achieved simultaneously in some critical applications to achieve the application requirements. However, this is challenging due to the conflicting nature of these KPIs. This article proposes a new service class of 6G as immersive, hyper reliable, and low-latency communication (IHRLLC), and introduces a potential network architecture to achieve the associated KPIs. Specifically, technologies such as ultra-massive multiple-input multiple-output (umMIMO)-aided terahertz (THz) communications, and reconfigurable intelligent surfaces (RIS) are viewed as the key enablers for achieving immersive data rate and hyper reliability. Given the computational complexity involved in employing these technologies, and the challenges encountered in designing real-time algorithms for efficient resource allocation and management strategies as well as dynamic beamforming and tracking techniques, we also propose the involvement of other potential enabling technologies such as non-terrestrial networks (NTN), learn-to-optimize (L2O) and generative-AI (GenAI) technologies, quantum computing, and network digital twin (NDT) for limiting the latency.
@@ I Introduction
Recently, the fifth-generation (5G) networks have become operational in many countries worldwide, while researchers are working on studies related to the sixth-generation (6G) wireless networks. These studies are inspired by potential use cases, requirements, and challenges that are yet to be addressed. The previous generations of wireless networks have already provided significant enhancements that enabled a vast range of novel applications compared to those supported by legacy wireless networks. Nevertheless, according to the International Telecommunication Union’s (ITU) vision for international mobile telecommunications 2030 (IMT-2030), 6G usage scenarios are envisioned as advanced versions of existing 5G scenarios, alongside entirely new use cases that support the deployment of innovative applications [1]. Besides, 6G wireless networks are expected to cater to future use cases and applications that require high mobility, immersive data rates, hyper-reliability, and low-latency simultaneously. This necessitates the introduction of significantly more advanced network architectures and enabling technologies. We refer to this type of service class as immersive, hyper reliable, and low-latency communication (IHRLLC), which is also the focus of this article.
@@
The current 5G wireless network already includes ultra-reliable and low latency communications (URLLC) as one of its service classes. The URLLC class targets mission-critical applications and a wide range of other applications, including industry automation, intelligent transportation, health care, and intelligent societies. Several enabling schemes and technologies have been developed to meet the stringent requirements of URLLC in terms of reliability and latency. These include grant-free access, redundant packet transmission, short packet transmission, edge computing, software-defined network (SDN), and network function virtualization (NFV) while utilizing both licensed and unlicensed bands, as well as coordinated multipoint (CoMP) techniques[2]. However, the critical use cases and applications that are envisioned for the 6G wireless networks demand unprecedented data rates combined with hyper-low latency and reliability simultaneously. In this context, URLLC alone might not be sufficient to achieve the requirements set for IHRLLC. Furthermore, it is essential to note that the IHRLLC requirements conflict with each other, making them challenging to achieve. Therefore, novel network architecture designs, along with innovative schemes and technologies, must be explored and investigated.
Recently, a few works have started investigating the role of URLLC in addressing use cases and applications that require high connectivity, mobility, and high data rates in 6G wireless networks [3, 4, 5]. In [3], the potential of terahertz (THz) frequency band is explored for offering high-rate and highly-reliable low-latency communications in wireless virtual reality (VR) applications. It is observed that the probability of line-of-sight (LoS) links in THz play a crucial role in characterizing the end-to-end (E2E) delay. In [4], three high data rate connectivity scenarios, including ubiquitous, deep, and holographic connectivity, are proposed, leveraging URLLC. Machine-learning-based solutions are proposed to achieve the stringent requirements of the high data rate and URLLC. In [5], a high mobility communication scenario is considered and the impact of the Doppler spread on URLLC for vehicular communication is explored. To this end, orthogonal time frequency space (OTFS) modulation and sensing-aided predictive beamforming schemes are exploited for reliable and low-latency communications, respectively.
@@
The aforementioned previous works explore individual use cases that require high data rates while adhering to the URLLC constraints. In this article, we present several important use cases for 6G wireless networks under the proposed IHRLLC service class. We then discuss potential enabling technologies and propose a unified approach to integrate these technologies to support the IHRLLC services. The main contributions of this paper can be summarized as follows
@@
  • We discuss a set of enabling technologies to support the IHRLLC class. Additionally, a network architecture is proposed that combines these technologies to achieve the goal of IHRLLC. Challenges and open issues pertaining to the proposed architecture are also discussed.
  • We present a numerical simulation-based experiment to validate the high data rate via ultra-massive multiple-input multiple-output (umMIMO)-aided THz communication.
@@ II Requirements, Use Cases, and Challenges
As 6G wireless networks are expected to encompass a broad spectrum of novel use cases, applications, and services, more than one KPI must be met simultaneously to achieve acceptable levels of quality of experience (QoE). In this paper, we focus on some specific use cases that require the simultaneous satisfaction of three KPIs, namely, immersive data rates, hyper reliability, and low latency. The exact values of these KPIs vary depending on service expectations of the use cases. Certain use cases prioritize specific KPIs or demand stricter thresholds on a particular KPI to guarantee the desired performance. In the following section, we discuss a novel service class for 6G which we term the IHRLLC class, and discuss some key use cases and associated challenges.
@@ II-A Immersive Communication
Immersive communication (IC) will be one of the key service classes of the 6G wireless networks. It will introduce interactive services such as extended reality (XR), holographic communication, cloud gaming, high-definition video (4K and 8K) streaming, and immersive event experiences over cellular networks. These services demand a large volume of data and require immersive throughput values. For instance, a live VR concert, streaming to thousands of participants simultaneously, requires a wireless network capable of handling substantial data rates without compromising the user’s QoE. According to the ITU IMT-2030 vision, the envisioned throughput for IC peaks at $1$ Tbps, with the user-experienced data rate of at least $300$ Mbps[1], which far exceeds the capabilities of IMT-2020.
@@ II-B Hyper reliable and low-latency communication (HRLLC)
The HRLLC class, an extension of URLLC in 5G, will be another core service class in 6G alongside other classes. Latency in wireless networks is influenced by various components of the architecture. Key factors include network management strategies in the core network, signal processing schemes such as modulation and coding in the E2E parts of the network, and the propagation delay of the signal within the network. Achieving hyper low-latency values of $0.1$ to $1$ ms, as specified by IMT-2030, without ensuring a hyper-reliable network, renders 6G use cases infeasible. Specifically, for applications such as autonomous vehicles and industrial automation, the network must simultaneously meet the reliability standard of $10^{-5}\sim 10^{-7}$ while achieving hyper low-latency, thus enabling the HRLLC usage scenario.
@@ II-C Immersive, hyper reliable, low-latency communication (IHRLLC)
The two aforementioned service classes have been studied separately, each with associated KPIs. Nevertheless, a wide range of use cases require simultaneous fulfillment of a combined set of three KPIs (i.e., immersive data rate, hyper reliability and low latency) for an acceptable level of QoE. Hence, we propose IHRLLC as a new service class for 6G wireless networks. Below are some of the key applications under this service class.
@@ II-C1 Mobile tele-surgery
Mobile tele-surgery in an ambulance racing through the city with a patient needing an urgent surgical operation is a representative use case for IHRLLC. The surgeon, located in a hospital far away from the ambulance, is required to perform the critical procedure remotely to save the patient’s life as they are being rushed to the hospital. This situation requires not only IC to transmit substantial data volumes but also necessitates absolute reliability and hyper low-latency to ensure the safety and efficacy of the medical procedure.
@@ II-C2 Multisensory XR
XR refers to various types of reality, including augmented reality (AR), mixed reality (MR), and virtual reality (VR), encompassing all real-and-virtual combined environments and human-machine interactions with the assistance of handheld and wearable end-user devices. Multisensory XR applications have unique requirements, including bounded low latency and high uplink and downlink bandwidth. These are challenging requirements due to scale and mobility at various speeds (e.g., walking, driving, fast trains) as well as coverage and handovers. The 3rd Generation Partnership Project (3GPP) Release 18 promises to significantly improve support for higher data rates and capacity while also introducing power and latency reduction techniques to enable more immersive AR experiences.
@@ II-C3 High-frequency trading
In particular, high-frequency trading (HFT) is a trading method that utilizes powerful computer programs to execute a large number of orders in a very short period. This approach aims to avoid the high risks associated with fast-changing trading prices[6]. HFT requires networks that are capable of meeting the required set of KPIs with a high degree of certainty.
@@ II-C5 Advance air mobility
Advance air mobility (AAM) is a concept that enables customers to access on-demand air transportation through multimodal transport networks. One of the key use cases of AAM is urban air mobility (UAM), which allows vehicles to either be autonomously controlled or remotely piloted while operating beyond the visual LoS from the ground. UAM vehicles typically have multiple sensors, such as light detection and ranging (lidar), radar, and cameras. These sensors can be utilized for advanced driver assistance systems and can also transmit real-time immersive data with hyper reliability and low-latency to a ground pilot for remote operation.
Pursuing a combined set of conflicting KPIs leads to inherent difficulties, as most existing wireless networks normally focus on achieving a specific KPI at the expense of others. Balancing these KPIs in 6G wireless networks to meet the diverse requirements of various applications of IHRLLC remains a significant challenge, and addressing them necessitates innovative solutions. In the following section, we overview a set of potential enabling technologies that we can depend upon and describe how properly aggregating these technologies overcomes the challenges that arise in simultaneously achieving conflicting KPIs.
@@
(a) CDF of per user data rate (Gbps).
(b) Packet Error Rate vs. SNR.
Figure 2: User data rate and PER in an umMIMO system working at $850$ GHz carrier frequency with $200$ MHz bandwidth.
@@ III Envisioned Enabling Technologies for IHRLLC
Several solutions and technologies are proposed to improve the throughput, latency, and reliability of wireless networks. These technologies span different parts of the wireless networks to ensure that the overall requirements are met. However, techniques used to achieve one specific KPI may interfere with achieving other KPIs. For example, utilizing higher frequency bands significantly improves data rates but limits the coverage area due to severe shadowing, deep fading, and molecular absorption, that result in decreased reliability. In this section, we provide an overview of our envisioned IHRLLC enabling technologies and their corresponding advantages and challenges. We also discuss how the combination of various technologies, across the different network layers, supports the concurrent fulfilment of multiple KPIs.
@@
Migrating to higher frequency bands has been the solution to access wider bandwidth and, hence, higher data rates for users. The millimeter wave (mmWave) spectrum, which includes frequencies above $24$ GHz, was utilized in the 5G wireless networks to provide access to a considerable bandwidth compared to the previous mobile network generations. However, the huge capacity demand of 6G applications necessitates the provisioning of extremely high bandwidths, which can only be achieved through THz communications. The THz spectrum, in particular, spans the frequencies in the $0.1$ to $3$ THz range. Extremely high available bandwidth at this frequency range makes it a promising technology to achieve immersive communication in 6G. Nevertheless, operating in the THz frequency band exposes new challenges to the wireless networks [7]. The most significant of these challenges include higher free space path loss (FSPL) and molecular absorption, channel modeling, and the design of hardware which includes radio frequency (RF) devices, modulators, and antennas. Therefore, devising innovative solutions to address the aforementioned challenges is crucial for leveraging the potential of THz frequencies in 6G wireless networks. To overcome the short distance communication range and severe blockage of the THz frequency band, antenna array architectures, e.g., umMIMO and RIS, can be integrated into the radio access networks to improve the beam gain toward users and maintain LoS connection with the base station, respectively, see Fig. 1.
@@
Higher Frequency bands, such as mmWave and THz, enable the use of antenna designs that can accommodate a significantly greater number of antenna elements. Therefore, implementing the umMIMO technology and enhanced beamforming techniques creates high-gain narrow beams tailored to individual users. These beams help reduce the substantial FSPL in THz communications[8]. Fig. 2 depicts the cumulative distribution function (CDF) of the data rate and the packet error rate (PER) performance for varying values of signal-to-noise ratio (SNR) in a wireless system consisting of one base station, with a cell radius of $1000$ m, equipped with an umMIMO antenna serving one user. The system works at the center frequency of $850$ GHz, and the user is allocated a bandwidth of $200$ MHz. It can be observed that the data rate and PER values of the user are within the desired range to facilitate a reliable and immersive connection. Nevertheless, achieving such high data rate and low PER values necessitates designing efficient advanced beamforming algorithms.
Beamforming via umMIMO architecture in the THz frequency band presents several challenges. Very narrow pencil beams require accurate knowledge of the channel state information (CSI), accurate angle of departure of the transmitters, and the angle of arrival of receivers. Furthermore, since beamforming is vulnerable to mobility, the difficulty lies in continuous beam realignment in a short time. Furthermore, a large number of RF power amplifier chains are needed to perform effective beamforming, thus increasing cost and power consumption. Alternatively, an energy-efficient holographic MIMO system appears to be a potential solution.

### v2

@@ Abstract
The transition towards the sixth-generation (6G) wireless telecommunications networks introduces significant challenges for researchers and industry stakeholders. The 6G technology aims to enhance existing usage scenarios through supporting innovative applications that require stringent key performance indicators (KPIs). In some critical use cases of 6G, multiple KPIs, including immersive throughput, with an envisioned peak data rate of $\mathbf{1}$ Tbps, hyper-reliability, in the range of $\mathbf{10^{-5}}$ to $\mathbf{10^{-7}}$, and hyper low-latency, between $\mathbf{0.1}$ and $\mathbf{1}$ ms, must be achieved simultaneously to deliver the expected service experience. However, this is challenging due to the conflicting nature of these KPIs. This article proposes a new service class of 6G as immersive, hyper reliable, and low-latency communication (IHRLLC), and introduces a potential network architecture to achieve the associated KPIs. Specifically, enhanced technologies, such as ultra-massive multiple-input multiple-output (umMIMO)-aided terahertz (THz) communications, reconfigurable intelligent surfaces (RIS), and non-terrestrial networks (NTN), are viewed as the key enablers for achieving immersive data rates and hyper reliability. Given the computational complexity involved in employing these technologies, we propose mathematical and computational enabling technologies, such as learn-to-optimize (L2O), generative-AI (GenAI), quantum computing, and network digital twin (NDT), to complement the proposed architecture and optimize the latency.
@@ I Introduction
Recently, the fifth-generation (5G) networks have become operational in many countries worldwide, while researchers are working on studies related to the sixth-generation (6G) wireless networks. These studies are inspired by potential use cases, requirements, and unresolved challenges. The previous generations of wireless networks have already provided significant enhancements that enabled a vast range of novel applications compared to those supported by legacy networks. Nevertheless, according to the International Telecommunication Union’s (ITU) vision for international mobile telecommunications 2030 (IMT-2030), 6G usage scenarios are envisioned as enhanced versions of existing 5G scenarios, alongside entirely new use cases that support the deployment of innovative applications [1]. Additionally, 6G wireless networks are expected to cater to future use cases and applications that require the concurrent fulfillment of high mobility, immersive data rates, hyper-reliability, and low-latency. This necessitates the introduction of significantly more advanced network architectures and enabling technologies. We refer to this type of service class as immersive, hyper reliable, and low-latency communication (IHRLLC), which is also the focus of this article.
@@
The current 5G wireless network already includes ultra-reliable and low latency communications (URLLC) as one of its service classes. The URLLC class targets a wide range of use cases, including industry automation, intelligent transportation, health care, and intelligent societies. Several enabling schemes and technologies have been proposed to meet the stringent requirements of URLLC in terms of reliability and latency [2]. These include grant-free access, spatial diversity, short packet transmission, slicing, edge caching and computing, etc. However, the critical use cases and applications that are envisioned for the 6G wireless networks demand unprecedented data rates combined with hyper-low latency and reliability simultaneously. In this context, URLLC alone might not be sufficient to achieve the requirements set for IHRLLC. Furthermore, it is essential to note that the IHRLLC requirements conflict with each other, posing significant challenges. Therefore, novel network architecture designs, along with innovative schemes and technologies, must be explored and investigated.
@@
Recently, a few works have started investigating the role of URLLC in addressing use cases and applications that require high connectivity, mobility, and high data rates in 6G wireless networks [3, 4, 5]. In [3], the potential of the terahertz (THz) frequency band is explored for offering high-rate and highly-reliable low-latency communications in wireless virtual reality (VR) applications. In [4], three connectivity scenarios requiring high data rates, namely ubiquitous, deep, and holographic connectivity, are proposed, leveraging URLLC. In [5], a high mobility communication scenario is considered, and the impact of the Doppler spread on URLLC for vehicular communication is explored. The aforementioned works explore individual use cases that require high data rates while adhering to the URLLC constraints. In this article, we present several critical use cases for 6G wireless networks under the proposed IHRLLC service class. Additionally, we discuss potential enabling technologies and propose a unified approach to integrate these technologies to support the IHRLLC services. The main contributions of this paper can be summarized as follows:
@@
  • We discuss a set of enabling technologies to support the IHRLLC class. Additionally, a network architecture is proposed which combines these technologies to achieve the goal of IHRLLC. Challenges and open issues pertaining to the proposed architecture are also discussed.
  • We present a numerical simulation-based experiment to validate the high throughout via ultra-massive multiple-input multiple-output (umMIMO)-aided THz communication.
@@ II Requirements, Use Cases, and Challenges
As 6G wireless networks are expected to encompass a broad spectrum of novel use cases, applications, and services, more than one KPI must be met simultaneously to achieve acceptable levels of quality of experience (QoE). Considering the end-user experience, QoE can be defined as “the overall acceptability of an application or service, as perceived subjectively by the end-user” [6]. In the following section, we discuss a novel service class for 6G which we term the IHRLLC class, and discuss some key use cases and associated challenges. Specifically, we focus on use cases that require the simultaneous satisfaction of three KPIs, namely, immersive data rates, hyper reliability, and low-latency.
@@ II-A Immersive Communication
Immersive communication (IC) will be one of the key service classes of the 6G wireless networks. It will introduce interactive services, such as extended reality (XR), holographic communication, cloud gaming, high-definition video (4K and 8K) streaming, and immersive event experiences over cellular networks. These services demand a large volume of data and require immersive throughput values. For instance, a live VR concert, streaming to thousands of participants simultaneously, requires a wireless network capable of handling substantial data rates without compromising the user’s QoE. According to the ITU IMT-2030 vision, the envisioned throughput for IC peaks at $1$ Tbps, with the user-experienced data rate of at least $300$ Mbps[1], which far exceeds the capabilities of IMT-2020.
@@ II-B Hyper Reliable and Low-Latency Communication (HRLLC)
The HRLLC class, an extension of URLLC in 5G, will be another core service class in 6G. Latency in wireless networks is influenced by various components of the architecture. Key factors include network management strategies in the core network, signal processing schemes, such as modulation and coding in the end-to-end (E2E) parts of the network, and the propagation delay of the signal within the network. Achieving hyper low-latency values of $0.1$ to $1$ ms, as specified by IMT-2030, without ensuring a hyper-reliable network, renders 6G use cases infeasible. Specifically, for applications, such as autonomous vehicles and industrial automation, the network must simultaneously meet the reliability standard of ${10^{-5}}$ to ${10^{-7}}$ while achieving hyper low-latency, thus enabling the HRLLC.
@@ II-C Immersive, Hyper Reliable, Low-Latency Communication (IHRLLC)
The two aforementioned service classes have been studied separately, each with its own associated KPIs. Nevertheless, a wide range of use cases requires simultaneous fulfillment of the three KPIs (i.e., immersive data rate, hyper reliability, and low-latency) to achieve an acceptable level of QoE. Hence, we propose IHRLLC as a new service class for 6G wireless networks. The IHRLLC service class must ensure that the three KPIs are consistently met, with zero tolerance for service interruptions. Below are some of the key applications under this service class.
@@ II-C1 Mobile tele-surgery
Mobile tele-surgery in an ambulance racing through the city with a patient needing an urgent surgical operation is a representative use case for IHRLLC. This situation requires not only IC for transmitting substantial data volumes but also absolute reliability and hyper low latency to ensure the safety and efficacy of the medical procedure.
@@ II-C2 Multisensory XR
XR refers to various types of reality, including augmented reality, mixed reality, and VR. Multisensory XR applications have unique requirements, including bounded low latency and high uplink and downlink bandwidth. The 3rd Generation Partnership Project (3GPP) Release 18 promises to significantly improve support for higher data rates and capacity while also introducing power and latency reduction techniques to enable more immersive XR experiences.
@@ II-C3 High-frequency trading
High-frequency trading (HFT) is a trading method that utilizes powerful computer programs to execute a large number of orders in a very short period. This approach aims to avoid the high risks associated with the fast-changing trading prices. HFT requires networks that are capable of meeting the required set of KPIs with a high degree of certainty.
@@ II-C5 Advance air mobility
Advance air mobility (AAM) enables customers to access on-demand air transportation through multimodal transport networks. One of the key use cases of AAM is urban air mobility (UAM), which allows vehicles to either be autonomously controlled or remotely piloted while operating beyond the visual line-of-sight (LoS) from the ground. The UAM optimal performance requires real-time immersive data transmission with hyper reliability and low-latency to a ground pilot for remote operation.
Pursuing a combined set of conflicting KPIs leads to inherent difficulties, as most existing wireless networks usually focus on achieving a specific KPI at the expense of others. Balancing these KPIs in 6G wireless networks to meet the diverse requirements of IHRLLC remains a significant challenge, and addressing them necessitates innovative solutions. In the following section, we overview a set of potential enabling technologies and describe how properly aggregating these technologies overcomes the challenges that arise in simultaneously achieving conflicting KPIs.
@@ III Envisioned Enabling Technologies for IHRLLC
Several solutions and technologies are proposed to improve the throughput, latency, and reliability of wireless networks. These technologies span different parts of the wireless networks to ensure that the overall requirements are met. However, techniques used to achieve one specific KPI may interfere with achieving other KPIs. For instance, utilizing higher frequency bands significantly improves data rates but limits the coverage area due to severe shadowing, high path loss, and molecular absorption, ultimately affecting reliability. Similarly, techniques like edge computing, which prioritize low latency, may struggle to scale effectively in scenarios requiring hyper-reliable communication. In this section, we provide an overview of our envisioned IHRLLC enabling technologies, highlighting their advantages and challenges. To address the trade-offs between different KPIs, we propose leveraging complementary technologies that work together to mitigate these challenges. Additionally, we discuss how integrating diverse technologies across various network layers can facilitate the simultaneous achievement of multiple KPIs.
@@
Migrating to higher frequency bands has been the solution to access wider bandwidth, thereby enabling higher data rates for users. The millimeter wave (mmWave) spectrum, which encompasses frequencies above $24$ GHz, was utilized in the 5G wireless networks to provide access to a considerable bandwidth compared to earlier mobile network generations. However, the immense capacity demands of 6G applications necessitate access to even broader bandwidths, achievable only through THz communication. The THz spectrum spans frequencies from $0.1$ to $3$ THz. Extremely high available bandwidth at this frequency range makes it a promising technology to achieve immersive communication in 6G. Nevertheless, operating in the THz frequency band exposes new challenges to the wireless networks [7]. The most significant challenges include higher free space path loss (FSPL), molecular absorption, channel modeling, and the design of hardware components, such as radio frequency (RF) devices, modulators, and antennas. Therefore, innovative solutions are essential to address the aforementioned challenges and fully harness the potential of THz frequencies in 6G wireless networks. For instance, to overcome the short communication range and severe blockage issues associated with the THz band, advanced antenna array architectures, e.g., umMIMO and reconfigurable intelligent surfaces (RIS), can be integrated into radio access networks to improve the beam gain toward users and help maintain LoS connection between the base stations and users, respectively (Fig. 1).
Figure 2: Per user throughput and BER versus the number of users for umMIMO-enabled THz communication.
@@
Higher frequency bands, such as mmWave and THz, enable the use of umMIMO antenna designs that can accommodate a significantly greater number of antenna elements, compared to the traditional MIMO architectures. As a result of implementing the umMIMO technology and enhanced beamforming techniques, high-gain narrow antenna beams can be tailored to individual users. These beams help overcome the substantial FSPL and short-range communication in THz communications [8]. Fig. 2 plots the throughput and bit error rate (BER) performance under different numbers of antenna elements versus the number of users in the umMIMO-enabled THz communication system. The network consists of a base station and up to $20$ IHRLLC users in a cell of radius $500$ meters, operating at a center frequency of $850$ GHz with bandwidth of $2$ GHz. The network resources are equally and orthogonally divided among users to eliminate interference. It is observed that the throughput and BER values remain within acceptable ranges, supporting reliable and immersive communication. The performance improves with higher-order umMIMO antennas; for instance, upgrading from a $512\times 512$ antenna array to a $1024\times 1024$ array yields slight throughput improvement. This moderate gain is influenced by the logarithmic nature of the relationship, but further enhancements can be expected with even higher-order umMIMO arrays. Nonetheless, achieving such high data rates and low BER for a large number of users with limited resources highlights the necessity of advanced beamforming algorithms and strong error control coding techniques to optimize performance.
umMIMO beamforming in the THz frequency band presents several challenges, including the need for precise channel state information (CSI), accurate angle of departure of the transmitters, and the angle of arrival of receivers. Additionally, beamforming is highly sensitive to mobility, necessitating rapid and continuous beam realignment. Moreover, a large number of RF power amplifier chains are needed to perform effective beamforming, thus increasing the cost and power consumption. Alternatively, an energy-efficient holographic MIMO system appears to be a potential solution.